if cot of theta is 4/5, and in quadrant 3, what is the sin of theta?
Answer by lwsshak3(11628)

if cot of theta is 4/5, and in quadrant 3, what is the sin of theta?
cot theta=4/5=adj side/opp side
hypotenuse
=√(4^2+5^2)
=√(16+25)
=√41
sin theta=opp side/hypotenuse=-5/√41 (sin<0 in quadrant 3)